# A Data-to-Product Multimodal Conceptual Framework to Achieve Automated Software Evolution for Context-rich Intelligent Applications

### v1

@@ 3.1 Context Sources and Data Modalities
Context sources for the software design, development and maintenance encompass various activities of any phases of the software engineering life cycle, as well as internal and historical data from intelligent applications, and correlated enviroments. The underlying philosophy behind incorporating such a broad spectrum of context sources is that the more related details are collected, the better decisions can be made in generating the final software products. This assumption is based on a hypothesis that the learning methods can effectively leverage the diverse data pool. Contextual data derived from applications includes users’ habits and other historical information, alongside sensor data from the surrounding environment, which can help to capture whether a user is correlated with a specific state, such as weather temperature, body movement, noise level, and more.
@@ 3.2 Multimodal Learning
The definitions of mulitmodalities are comprehensively studied by Letitia et al. [35]. Three types of definitions are discussed in their work, including human-centered (relies on human perceptual experience such as hearing and seeing), machine-centered (how is information represented), and ML task-related (inputs and outputs are represented differently or are composed of distinct types of atomic units of information).
There are a bunch of techniques in multimodal learning. To better describe this component, we use a categorization method by dividing them into white-box, black-box, and gray-box ML models. Meanwhile, multimodal learning also considers how to preprocess data and the fusion of the data.
So this component in our framework consists of the subcomponents of representing data processing tasks and machine learing methods that can yields different inputs or outputs or distinct types of atomic units of informaiton.
Among those methodologies, we particularly categorize them into three types of techniques: black-box learning, white-box learning, and grey-box learning.
Black-box, white-box and grey-box, refer to the last section.
Not always different types, but also even the text modality, it can have different sources, different data collection of different types, and can be multimodalities in text format. For example, multimodalities used in this research.
@@ 3.3 Key Features and Products
Key features are essential components of the learning process, serving as the building blocks for the final products. However, the ultimate objective remains the creation of useful products. This perspective underscores a shift towards a data-to-product approach, emphasizing the transformation of data into tangible outcomes rather than focusing solely on feature generation. Consequently, the automated software evolution process must encompass a decision-making loop to address this aspect effectively. The conceptual framework outlined here not only delves into the intricacies of various domains and subproblems but also explores methods of integrating them to create a more comprehensive framework.
The paper delves into the realm of dynamic code generation [26], particularly focusing on competition-level code generation with AlphaCode.
These second-order effects are often observed due to the dynamic nature of human learning systems. Human users continually refine their models, beliefs, and expectations through trial and error, resulting in a learning process akin to reinforcement learning. In some cases, the learning dynamics may exhibit positive feedback loops, potentially amplifying smaller effects over extended periods (Sutton and Barto, 1998).
The technologies examined in this study aim to influence human user behavior towards specific objectives, such as increasing the time spent on a website. This goal-driven behavior has been previously discussed in literature [8]. It is imperative to critically analyze and regulate any technology with the potential to steer human behavior.
@@
The main tasks or products of software evolution are in the phases of SE, as shown in the rows of Figure 5. The reason why we developed the conceptual framework is to set up the concepts related to and their relations; as a result, we identify the products. We notice that some automation approaches cover two phases; some cover one, and some cover more. So, it is necessary to develop a model to describe the research coverage.
A Selective Sequential Scope Model (SSSM or 3S model) is developed for this purpose, which means that the model can be used to cover selected scopes (one or more phases). The reason it is sequential is because it is not practical to cover the environment and design without covering/considering the requirement phase by common sense. Although in the future, there could be a way to do that, but that will be another type of model. This model is for the sequential coverage of the scopes.
@@ 4.1 Tasks
In order to solve the change from data to product, it is necessary to consider another dimension, called data-logic-model, as shown in Figure 5.
The result of multimodal learning (in logic deduction) is a model or system that can perform tasks or make predictions based on a more comprehensive and holistic understanding of the input data. For example, in a multimodal model that processes both textual and visual data, the product could be improved performance in tasks like image captioning, sentiment analysis on multimedia content, or any application where combining information from different modalities provides a richer and more accurate representation of the underlying patterns or semantics.
Data preprocessing can be extracting initial features, or high level features, just as analogy from image processing. To learn some high level features, some initial ML algorithms can be applied already. Logic is the business logic, or the entity relationship logic. Model is the complex function to resolve the decision task given input data. The first one is for intepretability and sometimes the second one is a blackbox, although sometimes it can be a white box.
the logic deduction means the reasoning processes, including using LLM or other techniques such as knowledge-graph construction. finally, there is a system or model that can be used to represent the logic, knowledge graph can be learned.
Logic deduction can be represented in text information, or other model, or in the data model way, then, formal model can be a specific language implementation of the logics. Data processing include data modeling and preprocessing, and data fusion, then logic deduction will be using algorithms, including ML, AI, or mutlimodality, to generate the decisions, which can be another model described using ML model.
The formal Method will be a way to represent those generated ML models, or it is a visualization of the ML model; either way, it can be used to represent the result of learning from the data. Are there some examples of like (t1, p2) ? not sure, I need to make literature review.
@@ 4.2 Software Engineering (SE) Phases
The main body of the SE phases is from the classical waterfall model. However, we are adding phase zero as the start phase and phase seven as a flexible phase. Phase zero is the environment from which all the requirements and context data are based. Phase seven is named "PhasingOut" in Figure 5. However, it can be documenting, configuration, or other things.
Particularly, phasingOut stage taks care of how to store or delete the data with considering the privacy protection and other regulations, meanwhile, the data processed in this stage may preserve the business logic and database logics so that it can be used for future development.
@@ 4.3 Scope
In this model, we list two dimensions: tasks and SE phases. The scope of the 3S model means the intersections of the two dimensions covered by an ASEv research. The research can be analyzed with respect to its relationship to the two dimensions in order to decide which scope it relates to. For example, if a study explores using UML design to generate the code and the test cases in an automated way, we say it covers p2 to p4 (including p3) in the SE phase dimension. If it is using data processing and logic deduction, it also covers t0 to t1 in the task dimension.
Some notations are defined to describe the scope in a simplified way. Firstly, we use a pair to represent the coverage in two dimensions. Secondly, we use >and <to describe the direction of the coverage. If both are used, it means that the study allows the affection of phases in two ways, e.g., P3 <>P4, the data model learned from the development phase can generate the test cases, and meanwhile, the model generated from the test phases can help with the code development. If there is no continuous coverage, then a comma is used to include indexes of phases or tasks that are not contiguous to each other. The arrows shown in Figure 5 are used to denote scopes. The type with a wider area covering multiple columns considers two or more tasks simultaneously. The type that covers only one column involves one task.
@@ 5 Apply the Framework and 3S Model to ASEv-related Research
This section presents some endeavors that targets ASEv or can help with ASEv. We apply the data-to-model conceptual framework to analyze those studies and categorize them using the 3S model.
@@ 5.1 Agent-based Software Development
ChatDev [39], a virtual chat-powered company for software development, brings together "software agents" from diverse social identities, including chief officers, professional programmers, test engineers, and art designers. "collaborative chatting" among those agents will take place when presented with a human client’s request. The method of ChatDev enables automatic crafting comprehensive software solutions that encompass source codes, environment dependencies, and user manuals.
ChatDev employs the widely adopted waterfall model and divide the software development process into four distinct phases: designing, coding, testing, and documenting. Particularly, a chat chain is designed to facilitate breaking down each phase into atomic subtasks, allowing for proposing and validating solutions through context-aware communication.
Their work with agents depends on AI tools such as ChatGPT to generate code and evaluate the code, across the stages of designing coding, testing and documenting. Since documenting is not listed in the 3S model, we just consider the first three stages. To add documenting or other phases to the model, a customized extension can be made to the 3S model. Thus, according to the scope definition in the 3S model, we can use the pair (t2, p2>4) to represent the scope of this research. Similarly, we generated the scopes of the rest research as shown in Table 1.
@@
Agent-Based
ChatDev [39]
chat transcript, LLM
(t2, p2>4)
Machine learning
DeepFix [16]
C programs
(t0>2, p4)
CodeBERT[12]
source code
(t0>2, p4)
ArduCode [4]
souce code, textual metadata, hardware configuration lists
(t0>2, p1&3)
Genetic Programming
GenProg [24]
souce code, test cases
(t0>1, p3<>4)
Ontology-based
A lightweight semantic processing approach [22]
requirement specification, ontological elements
(t0>2, p1)
Stated-based
CSMEngine [51]
high level context data
(t0>2, p0>2)
5.2 Machine learning
Fixing bugs is an essential activity in the maintenance of software. Bugs can be identified in the testing phase, or can be reported through analyzing the failure information of a running system. No matter how the bugs are identified, they are supposed to be corrected. In this study, bug fixing is considered to be in the development and testing phases in the waterfall model.
DeepFix [16] is a method for fixing common C language errors by deep learning. They treat the problem of fixing a programming error as a sequence-to-sequence learning problem, which requires a program to be represented as a sequence. They constructed the dataset by collecting correct programs and erroneous programs. Through the mutation of correct programs, they can build training examples which consists of the mutated programs and the fix of them. Their trained model achieve a successful fix rate of 27%. By collecting data from code examples and design the deep learning neural network architecture for an attention based sequence-to-sequnce model, long term dependencies can be captured for utilizing the context in fixing the bugs.
CodeBERT[12] is a transformer-based neural architecture pre-trained on large corpus of source code. Ehsan et. al [30] apply it for automated program repair of Java simple bugs. Through experiments, they find the method can generates fix codes in 19-72% of the cases with different types of their datasets, which are exactly the same as the fix codes implemented by developers.
ArduCode [4] is a predictive framework for assisting the automatioin engineers in classifying automation code, finding similar code snippets and reasoning about the hardware selection of sensors and actuators to resolve the requirement challgenes from hardware and software integration. They use paragraph enmbedding in classification and achieve precision close to human annotation. Hardware recommendation are studied using autoencoder models. Thus, their work is categorized in the 3S model as working with the requirement and development.
@@ 5.5 State-based
CSMEngine [51] is proposed to automatically model and reason high-level context. Their study applied the methodology to an intelligent elevator system. CSMEngine is based on context state machine modeling methods [52] and CSM-H-R [50] is a framework as an extension of the core model, to facilitate the modeling of the context dynamism. CSMEngine works with high-level context information, which is the data fused from the environment or applications. By automatically modeling the collected context, it can reason the context to identify new rules the applications can adapt to or the design process can utilize.
@@
The research presented in this section is not based on a thorough survey. However, to the best of our knowledge, there is little research that covers more than two stages in the SE phase dimension. As more data will be collected and the application of multimodal learning to more specific tasks, we predict there will be research that can cover a larger scope so that a higher level of automation of the software evolution process can be achieved.

### v5

@@ 3.1 Context Sources and Data Modalities
Context sources for software design, development, and maintenance encompass various activities in any phase of the software engineering life cycle, as well as internal and historical data from intelligent applications and correlated environments. The underlying philosophy behind incorporating such a broad spectrum of context sources is that the more related details are collected, the better decisions can be made in generating the final software products. This assumption is based on a hypothesis that the learning methods can effectively leverage the diverse data pool. Contextual data derived from applications includes users’ habits and other historical information, alongside sensor data from the surrounding environment, which can help to capture whether a user is correlated with a specific state, such as weather temperature, body movement, noise level, and more.
@@ 3.2 Multimodal Learning
The definitions of mulitmodalities are comprehensively studied by Letitia et al. [49]. Three types of definitions are discussed in their work, including human-centered (relies on human perceptual experience such as hearing and seeing), machine-centered (how is information represented), and ML task-related (inputs and outputs are represented differently or are composed of distinct types of atomic units of information). According to the ML task-related definition, it is multimodal learning even if only text data is involved and if the input and output formats are different or it is handling different units of data.
There are a bunch of techniques in multimodal learning. To better describe this component, a categorization method is particularly used by dividing them into white-box, black-box, and gray-box ML models. Firstly, they are terms that originated in the software engineering area. Secondly, the three categories can provide a complete coverage of the models. In each of the categories, specific methods are listed in the framework component, as shown in Figure 4. Meanwhile, multimodal learning also considers how to preprocess data and the fusion of the data. So, this component in our framework consists of the subcomponents representing data processing tasks and machine learning methods that work with different inputs or outputs or distinct types of atomic units of information.
@@ 3.3 Key Features and Products
Key features of ASEv, as the objectives of multimodal learning, are essential components of the learning process. They are functional modules for building final products, including new features, programs, systems, and their updates. After all, the ultimate objective remains to be creating useful products. This perspective underscores a shift towards a Data-to-Product approach (the connection from D to P in Figure 4), emphasizing the transformation of data into tangible outcomes rather than focusing solely on the evolution procedures. Consequently, the automated software evolution process must encompass a decision-making loop to address this aspect effectively, e.g., a new feature can be examined by a new phase of learning. Thus, the conceptual framework outlined here not only delves into the intricacies of various domains and subproblems but also explores the integration of them to fulfill the Data-to-Product view.
Feature generation and enhancement benefit the function designs and can be learned from the user’s habits and changes in the context of the environment. They are mapped with the software requirement phase. A feedback loop module is particularly necessary for the reason of the second-order effects, as discussed in [58], which can be observed due to the dynamic nature of human learning systems. When working with intelligent software agents, human users continually refine their models, beliefs, and expectations through trial and error, resulting in a learning process akin to reinforcement learning. In some cases, the learning dynamics may exhibit positive feedback loops, potentially amplifying smaller effects over extended periods.
Dynamic code generation deals with the software development phase. For example, the work [37] delves into the realm of dynamic code generation, particularly focusing on competition-level code generation with AlphaCode. System integration, as an important consideration of today’s complex system development, cannot be neglected in order to provide robust and fault-tolerant services. Automated bug detection and correction, as mapped to the phase of software testing, ensures the correctness and quality of the software. Configuration, deployment, and dependency management are other crucial factors that facilitate the automation of the overall software evolution process.
@@
The tasks in achieving ASEv through multimodality learning correlate with the SE phases, as shown in the rows of Figure 5. So do the products mentioned in the conceptual framework. One of the benefits of developing the conceptual framework is identifying products through analyzing the concepts related to ASEv and their relations. Each product is generated through an automation approach that can cover multiple phases of SE. Given some ASEv research working with generating a product mentioned in the ASEv conceptual framework, it is significant to have a way to clearly describe the research coverage in terms of the SE phases and tasks.
This section presents a Selective Sequential Scope Model (SSSM or 3S model) for categorizing research work that spans different scopes (one or more phases correlate with one or more tasks). It is sequential because it is practical to work on one phase based on its prior phase. As a counterexample, it is still not practical to directly connect the phases of environment and design without covering/considering the requirement phase. Although there could be ways to achieve that, which can happen in the future or beyond the author’s knowledge, another model could be used to describe those approaches. This model is solely for the sequential coverage of the scopes.
@@ 4.1 Tasks
In order to support the transition from data to product, it is necessary to consider another dimension, data-logic-model, as shown in Figure 5. Data processing can include data modeling, preprocessing, fusion, extracting initial features, or high-level features. Some ML algorithms can be applied initially to learn some high-level features. Logics can include business logic or entity-relationship logic. The logic deduction is to use multimodality learning to generate the decisions, knowledge, or other types of models described using ML models.
Logic deduction is the reasoning process, including utilizing LLM or other multimodality learning techniques. The result of multimodal learning is a model or system that can perform tasks or make predictions based on a more comprehensive and holistic understanding of the input data. For example, in a multimodality model that processes both textual and visual data, the product can have improved performance in tasks like image captioning [68], sentiment analysis on multimedia content [57], or any other applications where combining information from different modalities provides a richer and more accurate representation of the underlying patterns or semantics.
The formal model is a way to represent those generated ML models or reasoning systems, or it is a visualization of them; either way, it can be used to represent the result of learning from the data. Generally speaking, models here refer to complex functions that represent the products or resolve the decision task given input data. Notably, general-purpose LLMs are not formal models in this dimension and are treated as a logic deduction tool if included in a methodology.
@@ 4.2 Software Engineering (SE) Phases
The main body of the SE phases is from the classical waterfall model. However, a special phase zero is added as the start phase, and phase seven is added as a flexible phase. Phase zero is the environment from which all the requirements and context data are based. Phase seven is named "PhasingOut" in Figure 5. However, it can be documenting, configuration, or other things.
Particularly, the "phasing out" stage takes care of how to store or delete the data while considering privacy protection and other regulations. Meanwhile, the data processed in this stage may preserve the business logic and database logic so that it can be used for future development.
@@ 4.3 Scope
This 3S model is presented with two dimensions: tasks and SE phases. The scope of the 3S model means the intersections of the two dimensions covered by an ASEv research. The research can be analyzed concerning its relationship to the two dimensions in order to decide which scope it relates to. For example, if a study explores using UML design to generate the code and the test cases in an automated way, it can be viewed as covering p2 to p4 (including p3) in the SE phase dimension. If it is using data processing and logic deduction, it also covers t0 to t1 in the task dimension.
Some notations are defined to describe the scope in a simplified way. Firstly, a pair is used to represent the coverage in two dimensions. Secondly, >and <are used to describe the direction of the coverage. If both are used, it means that the study allows the affection of phases in two ways, e.g., P3 <>P4, the data model learned from the development phase can generate the test cases, and meanwhile, the model generated from the test phases can help with the code development. If there is no continuous coverage, then a comma is used to include indexes of phases or tasks that are not contiguous to each other. The arrows shown in Figure 5 are used to denote scopes. The type with a wider area covering multiple columns considers two or more tasks simultaneously. The type that covers only one column involves one task.
@@ 5 Apply the Framework and 3S Model to ASEv-related Research
This section presents some endeavors that target ASEv or can help with ASEv. The data-to-model conceptual framework is applied to analyze these studies, and the 3S model categorizes them. It is not an exhaustive search of the literature, and only some representative research is selected as a demonstration.
In a broad sense, deep learning falls under the category of machine learning, and LLMs are special deep learning models that are trained on large amounts of data. In this section, we differentiate them into two categories, machine learning and LLM-based methods, mainly due to LLMs being trained on significantly more extensive and more diverse data corpora.
@@ 5.1 Agent-based Software Development
ChatDev [53], a virtual chat-powered company for software development, brings together "software agents" from diverse social identities, including chief officers, professional programmers, test engineers, and art designers. "collaborative chatting" among those agents will take place when presented with a human client’s request. The method of ChatDev enables the automatic crafting of comprehensive software solutions that encompass source codes, environment dependencies, and user manuals.
ChatDev employs the widely adopted waterfall model and divides the software development process into four distinct phases: designing, coding, testing, and documenting. Particularly, a chat chain is designed to facilitate breaking down each phase into atomic subtasks, allowing for proposing and validating solutions through context-aware communication.
Their work with agents depends on AI tools such as ChatGPT to generate code and evaluate the code across the stages of designing, coding, testing, and documenting. Since documenting is not listed in the 3S model, only the first three stages are considered. For adding phases such as documenting to the model, a customized extension can be made to the 3S model. Thus, according to the scope definition in the 3S model, the pair (t2, p2>4) can be used to represent the scope of this research. Similarly, the scopes of other research mentioned in this section are generated and shown in Table 1.
@@
Agent-Based ChatDev [53] chat transcript, LLM (t2, p2>4)
Genetic Programming GenProg [35] souce code, test cases (t0>1, p3<>4)
Ontology-based A lightweight semantic processing approach [31] requirement specification, ontological elements (t0>2, p1)
Case-based ACAI [16] specification text, code template (t2, p3)
Stated-based CSMEngine [70] high level context data (t0>2, p0>2)
Machine learning DeepFix [24] C programs (t0>2, p4)
ArduCode [8] source code, textual metadata, hardware configuration lists (t0>2, p1&3)
LLM-based CodeBERT[43] source code (t0>2, p4)
Empirical studies on code generation using LLM-based tools[46, 40] prompts, modalities of LLMs (t1, p1&3)
AID-an LLM-powered test case generation framework [39] prompts, modalities of LLMs (t1, p4)
@@ 5.3 Ontology-based
5.4 Case-based
An Automated coding system - ACAI (Automated Coder using Artificial Intelligence) is proposed by Danilchenko et al. [16]. They treat code generation as a design problem, and the solutions are plans. Thus, ACAI is built with combining the techniques of routine design, template-based programming, and case-based reasoning. By accepting the objective and the specifications for how to achieve the objective, ACAI yields the result in the form of a working Java program. ACAI highly relies on tailored input templates and program templates, and more importantly, the user of ACAI should design the target code by applying the mechanisms the system provides. Thus, this system is categorized as working on the formal model in the phase of development without covering automated data processing and logic deduction.
@@ 5.5 State-based
CSMEngine [70] is proposed to automatically model and reason high-level context. Their study applied the methodology to an intelligent elevator system. CSMEngine is based on context state machine modeling methods [71] and CSM-H-R [69] is a framework as an extension of the core model, to facilitate the modeling of the context dynamism. CSMEngine works with high-level context information, which is the data fused from the environment or applications. After automatically modeling the collected context, it can reason the context to identify new rules the applications can adapt to or the design process can utilize.
5.6 Machine learning
Fixing bugs is an essential activity in the maintenance of software. Bugs can be identified in the testing phase or reported by analyzing the failure information of a running system. No matter how the bugs are identified, they are supposed to be corrected. In this study, bug fixing is considered to be in the development and testing phases of the waterfall model.
DeepFix [24] is a method for fixing common C language errors by deep learning. They treat the problem of fixing a programming error as a sequence-to-sequence learning problem, which requires a program to be represented as a sequence. They constructed the dataset by collecting correct programs and erroneous programs. Through the mutation of correct programs, they can build training examples that consist of the mutated programs and the fixes. Their trained model achieved a successful fix rate of 27%. By collecting data from code examples and designing the deep learning neural network architecture for an attention-based sequence-to-sequence model, long-term dependencies can be captured to utilize the context in fixing the bugs.
ArduCode [8] is a predictive framework for assisting automation engineers in classifying automation code, finding similar code snippets, and reasoning about the hardware selection of sensors and actuators to resolve the requirement challenges from hardware and software integration. They use paragraph embedding in classification and achieve precision that is close to human annotation. Hardware recommendations are studied using autoencoder models. Thus, their work is categorized in the 3S model as working with the requirement and development.
5.7 LLM-based
CodeBERT[19] is a transformer-based neural model that has been pre-trained on an extensive corpus of source code. Ehsan et al. [43] apply it for automated program repair of Java simple bugs. Through experiments, they found the method can generate fix codes in 19-72% of the cases with different types of their datasets, which are exactly the same as the fix codes implemented by developers.
Nejjar et al. [46] conducted an empirical study on code generation tasks using LLM-based tools; specifically, the code generation tasks include a matrix multiplication function, scripts for data analysis, and data visualization. They performed the study by providing prompts to the tools, including GPT3.5, GPT4, Bing Chat, Google Bard, and YouChat. Their report shows the quality of code w.r.t. correctness, efficiency, comprehensibility, and overall rating. Liu et al. [40] conducted an empirical study on generating software code for safety-critical domains using GPT-4. Compared with Nejjar’s work, this study applied more approaches for utilizing the LLM tool, including overall requirements, specific requirements, and augmented prompts. The approach they applied is working on requirements and code development. The LLMs they used are general-purpose models, not formal models specifically for achieving ASEv tasks. They use LLMs only as a logic deduction tool.
Liu et al. [39] explore an LLM-powered test case generation method. A framework called AID is proposed. The framework consists of three main stages: the generation of program variants, the generation of program inputs, and differential testing. AID first feeds problem descriptions, program descriptions, and programs under test into LLMs to generate program variants and test inputs; then a diversity-first differential testing is applied to create test cases. AID is considered to be working in the phase of testing and the task of logic deduction.
5.8 Summary
@@
The research presented in this section is not based on a comprehensive survey. However, to the best of the authors’ knowledge, little research covers more than two stages in the SE phase dimension. As more data is collected and multimodal learning is applied to more specific tasks, it is predicted that future research will cover a broader scope so that a higher level of automation of the software evolution process can be achieved.
6 Discussion
This section presents an anticipated future trend and related prospective exploration directions and discusses some interesting topics that are closely relevant to the conceptual framework but were not discussed thoroughly in the main body due to the article’s focus on multimodality and data-to-product process. Open Research Questions (ORQ) are specially identified for advancing the work of the proposed ASEv conceptual framework and the 3S model.
6.1 Future Trends
A trend is anticipated in research regarding leveraging data across multiple software engineering phases. Several prospective avenues for exploration are envisioned as follows:
  1. 1.
    Initiating tasks, such as code generation, from raw environmental data or other types of contextual data.
  2. 2.
    Implementing logic adaptation strategies to dynamically adjust software behavior based on evolving requirements.
  3. 3.
    Harnessing Artificial General Intelligence [33] to map and synthesize code logic, followed by iterative optimization and refinement processes.
6.2 Full Automation and Semi-automation
All the aforementioned ASEv features could be approached through two methods: full automation (with minimum exterior intervention, e.g., initial configuration) and semi-automation (with significant exterior assistance, e.g., needing help in data generation, logic deduction). When simple tasks can be initiated through automated means, complex tasks necessitate an initial semi-automated approach, with the potential to transition gradually to full automation, for the initial stages of achieving full automation may pose technical challenges. This study presents a conceptual framework to facilitate the full automation of software evolution. Identifying tasks suitable for full automation and those requiring a semi-automated approach can help to apply a gradual progression toward automation as data thresholds are met and accumulated.
6.3 X-Box Learning
Explainability does not come for free. In order to make the model and learning process more understandable, extra efforts should be made to combine white-box learning with black-box learning to represent the knowledge in a way that human beings can comprehend. Although black-box learning models may be more efficient in time when dealing with large datasets and more accurate, as shown in many studies in the past decades, it does not mean they are always more accurate. The work of Emmanuel et al. [52] suggests white-box models or gray-box models can outperform black-box models, especially when the data amount is small. So, the real performance of various ASEv tasks involved needs to be investigated in practice.
6.4 Practical Managerial Significance
Management is a crucial facet in software engineering practices and theories [23, 32, 61], including project management, people management, progress management, quality management, risk management, and so on. The ’manager side’ can be human participants and management tools, while the ’managed side’ consists of human participants, software, and other control points. Practical Managerial Significance (PMS) can be a significant concern, as the ASEv process can affect the human participants and the software on both sides. Although this article primarily focuses on technical aspects and thus does not delve deeply into management concepts within the framework or the 3S model, exploring the integration of management practices in the framework and discussing the ASEv’s PMS remains an interesting direction for future research.
6.5 Data Amount and Privacy
The prerequisites of multimodality learning rely on the availability of sufficient data; if the data does not meet a threshold, effective training sometimes becomes impossible. Privacy considerations often present obstacles to data acquisition, and the concerns should be thoroughly discussed and resolved. For some types of intelligent systems, if they are for public usage and have few privacy concerns, obtaining data from those intelligent systems’ context could be more accessible. Nevertheless, it might still be hard to get the data from the software development cycles to protect the privacy or other ethical concerns of the software development participants. Possible strategies can be leveraging data from different software types and conducting cross-domain research to mitigate data scarcity within specific software development cycles.
6.6 Other Ethical Concerns
There is a term in sociology called Technological Determinism (TD) [25], which suggests that technology is not just a tool for human users to achieve goals, but rather it controls and shapes human behaviors. This subsection is not to support or discuss TD, but among many influences that technologies can bring to human society, persuasion can particularly be an ethical concern of ASEv and the framework due to the applications of black-box models such as LLMs, which exhibit the traits of persuasion [9] and the possibility of being persuaded [72]. In the analysis of the interaction between intelligent software agents and human users [7], technologies to influence human user behavior towards specific objectives are examined, such as increasing the time spent on a website. This goal-driven behavior has been previously discussed in the literature [14]. It is imperative to critically analyze and regulate any technology with the potential to steer human behavior in the processes and products of ASEv.
6.7 Limitations and Open Research Questions
This conceptual framework considers the software engineering phases as the context sources and as one dimension in the 3S model. However, it’s important to note that different methodologies of a specific phase can influence the workflow in a nuanced way, a factor that needs to be carefully considered. For instance, as a crucial consideration in the design phase and from the perspective of comparing technical-driven and tailored-usability-driven, system design methodologies can include hard systems, soft systems [11], and socio-technical design [4] as a relatively balanced approach. Soft systems take into account human activity dynamics and their tailored needs [6], while hard systems focus more on techniques and engineering methods. Such distinctions can lead to different data or logic flows under the framework of ASEv. Thus, the below ORQs are proposed for the framework and the 3S model.
  1. 1.
    Extend the framework by considering various methodologies of SE phases. For example, in the design phase, the system design methodologies can be emphasized for a particular type of software.
  2. 2.
    The framework is a general one that treats software as a whole. Software in different domains may have varying needs in the process of ASEv, so another dimension to consider is the specific domains. Besides, domains can be a third axis for the 3S model.
  3. 3.
    Expand the framework to incorporate ethical concerns and social impacts, such as PMS, providing valuable insights for software management.